# Perfect Matching in Product Graphs and in their Random Subgraphs

### v2

@@ 1.2 Proofs outline
With these definitions and tools at hand, we can now describe how to proceed in showing that typically obstructions do not exist. Using 1.2 one can deduce a lower bound of the number of edges in $E(K_{i},K_{i}^{C})$. Note that for $U$ to be an obstruction, $E_{G_{p}}(K_{i},K_{i}^{C})\subseteq E(K_{i},U)$ for all $i\in[\ell]$. Since $G$ is $d$-regular, there are at most $du$ edges touching $U$. So in order to have an obstruction in $G_{p}$, many of the edges in $\bigcup_{i\in\ell}E(K_{i},K_{i}^{C})$ cannot be present in $G_{p}$. Utilising the enumeration given by Lemma 3.7 and this probabilistic bound, we show in Lemma 3.8 that whp there are no obstructions with $1\leq u\leq\frac{n}{d^{C^{3}/p}}$. Then, using the enumeration given by Lemma 3.2 and the aforementioned probabilistic bound, we show that whp there are no obstructions with $b$ not ‘too large’. For larger values of $b$, the aforementioned probabilistic bound no longer suffices. Thus, the final ingredient, similar in spirit to the approach of Bollobás in [4], will be Lemma 3.5, which allows to efficiently bound the number of choices for sets $B$ that have ‘bad expansion’, that is, the number of edges in $E(B,B^{C})$ is small. We note that this is the only place where we use the structure of $G$ as a product graph. Indeed, sets $B$ with bad expansion leave a large ‘fingerprint’ on some coordinates of the product, in particular, if $B$ intersects non-trivially with a projection $G(I)$ of $G$ onto some coordinates $I\subseteq[t]$, then $G(I)$ spans many edges of $E(B,B^{C})$.
@@
Finally, let us briefly comment on the difference between the proof here and the proof for the hypercube, given in [4]. Since the hypercube is a bipartite graph, it suffices there to apply Hall’s theorem, and in particular, one does not need to consider the entirety of the giant component, but only sets of size at most $\frac{n}{2}$ which expand badly. Hence, Harper’s isoperimetric inequality for the hypercube, together with a fingerprint argument given by Bollobás [4], allows one to show that whp there are no obstructions (bar isolated vertices) to a perfect matching. Here, since $G$ is not necessarily bipartite, one needs to consider Tutte-type obstructions, and in particular sets encompassing the entirety of the giant component. This, in turn, requires a much more delicate treatment, and in particular, to consider the structure of obstructions, and the more efficient enumerations given by Lemmas 3.2 and 3.7. This part of the analysis is the key novelty of this work.
@@ 2 Connectivity
We begin by showing that isolated vertices in $G_{p}$ are typically at distance at least two in $G$.
@@
Fix an edge $\{u,v\}\in E$. Since $G$ is $d$-regular, the total number of edges meeting $u$ and $v$ is $12d-$ 1. Thus, the probability that $u$ and $v$ are isolated vertices in $G_{p}$ is at most
$(1-p)^{2d-1}\leq(1-p)^{-1}n^{-2(1-\epsilon)}\leq n^{-2(1-2\epsilon)},$
where we used that $(1-p)^{d}=n^{-(1-\epsilon)}$. There are $\frac{dn}{2}$ edges to consider. Thus, by the union bound, the probability that two isolated vertices in $G_{p}$ are at distance one in $G$ is at most $\frac{dn}{2}n^{-2(1-2\epsilon)}=o(1)$. ∎
Let us further show the following ‘gap’ statement, which is an almost-immediate corollary of Theorem 1.2 and Lemma 1.3. We note that $d^{20}$ in the following lemma is chosen rather arbitrarily, and a much tighter bound can be obtained.
@@
Fix $k\in[2,d^{20}]$. Let us estimate the probability that there is a connected component $K$ of order $k$ in $G_{p}$. By Theorem 1.2, we have that $e(K,K^{C})\geq k(d-(C-1)\log_{2}k)\geq\frac{9kd}{10}$, where we used the fact that $k\leq d^{20}$. Let $\mathcal{T}_{k}$ be the set of trees of order $k$, and let $\mathcal{A}_{k}$ be the event that there exists a connected component of order $k$ in $G_{p}$. Thus, by the union bound and by Lemma 1.3,
$\mathbb{P}\left[\mathcal{A}_{k}\right]$ $\leq\sum_{T\in\mathcal{T}_{k}}(1-p)^{e(T,T^{C})}\leq n(ed)^{k-1}(1-p)^{\frac{9kd}{10}}\leq n^{1-(1-\epsilon)\frac{9k}{10}}(ed)^{k-1}\leq n^{1-(1-2\epsilon)\frac{9k}{10}}.$
@@
Whp there is a unique giant component in $G_{p}$ whose order is $n-o(n)$, and all the other components in $G_{p}$ are isolated vertices. Furthermore, every two isolated vertices are at distance at least two from each other.
@@
Let $p_{2}=\frac{1}{d^{2}}$, and let $p_{1}$ be such that $(1-p_{1})(1-p_{2})=1-p$. Note there exists $\epsilon^{\prime}>0$ such that $(1-p_{1})^{d}=n^{-(1-\epsilon^{\prime})}$, and thus we may apply Lemmas 2.1 and 2.2 on $G_{p_{1}}$. Furthermore, observe that $G_{p}$ has the same distribution as $G_{p_{1}}\cup G_{p_{2}}$.
We begin by percolating with probability $p_{1}$. Note that by Lemma 2.2, whp in $G_{p_{1}}$ there are only isolated vertices and components of order at least $d^{20}$. We begin by showing that after sprinkling with $p_{2}$, typically all the components of order at least $d^{20}$ merge. Let $W$ be the set of vertices in components of order at least $d^{20}$ in $G_{p_{1}}$. Let $A\sqcup B=W$ be a partition of $W$ which respects the components of $G_{p_{1}}$, that is, every component is contained either in $A$ or in $B$. We may assume that $|A|\leq|B|$ and let $a\coloneqq|A|$.
Let $A^{\prime}\coloneqq\left(A\cup N_{G}(A)\right)\setminus B$, where $N_{G}(X)$ is the external neighbourhood of $X$ in $G$, and let $B^{\prime}\coloneqq\left(B\cup N_{G}(B)\right)\setminus A^{\prime}$. By Lemma 2.1, whp every two isolated vertices in $G_{p_{1}}$ are at distance at least two from each other in $G$. We continue assuming this holds deterministically. Thus, every isolated vertex in $G_{p_{1}}$ is in the neighbourhood of $A$ or $B$, and we have that $V=A^{\prime}\sqcup B^{\prime}$, that is, $B^{\prime}=V\setminus A^{\prime}$. By Theorem 1.2, we have that $e(A^{\prime},B^{\prime})\geq\frac{a\log_{C}(n/a)}{C-1}\geq\frac{a\log_{C}2}{C-1}$, since $a\leq\frac{n}{2}$. We can extend these edges to $\frac{a\log_{C}2}{C-1}$ paths of length at most $3$ between $A$ and $B$. Every two such paths share at most two edges. Thus, we can trim the set of $\frac{a\log_{C}2}{C-1}$ paths to a set of $\frac{a\log_{C}2}{Cd^{2}}$ edge-disjoint paths between $A$ and $B$.
We now sprinkle with probability $p_{2}$. The probability none of these paths are in $G_{p_{2}}$ is thus at most $(1-p_{2}^{3})^{\frac{a\log_{C}2}{Cd^{2}}}\leq\exp\left\{-\frac{a}{d^{9}}\right\}$. For a union bound note that there are at most $\binom{n/d^{20}}{a/d^{20}}$ possibilities to choose the partition of $W$. Thus, the probability that there is a partition of $W$ such that there is no path between two components of order at least $d^{20}$ is at most
@@
Note that by Lemma 2.1, whp every two isolated vertices in $G_{p_{1}}$ are not connected by an edge of $G$. Hence, adding any edge touching an isolated vertex connects it to a component whose order is at least $d^{20}$ in $G_{p_{1}}$, and these components all merge whp. Hence, whp, there exists a unique connected component in $G_{p_{1}}$ whose order is at least $d^{20}$, and all the other components are isolated vertices, whose distance in $G$ is at least two. Finally, let $Y$ be the random variable counting the number of isolated vertices in $G_{p}$. Then, $\mathbb{E}[Y]=n(1-p)^{d}\leq n^{\epsilon}$. Thus, by Markov’s inequality, whp $Y\leq n^{2\epsilon}=o(n)$, completing the proof. ∎
@@ 3 Perfect matching
We begin with the proof of Theorem 1. Throughout the section, we assume divisibility by two whenever is necessary, noting that minor modifications will allow us to argue for the existence of nearly-perfect matching otherwise. Let us first show that, by Theorem 1.2, for ‘high enough’ dimension of $G$ every edge-cut of $G$ has at least $d$ edges.
@@
Suppose that $t\geq 5C\log_{2}C$. Then, we have that $G$ is $d$-edge-connected.
@@
It suffices to show that $e(S,S^{C})\geq d$ for all $S\subseteq V$ with $11\leq|S|\leq n-$ 1, and in fact by symmetry, for $S$ of size at most $\frac{n}{2}$. By 1.2,
$e(S,S^{C})\geq|S|(d-(C-1)\log_{2}|S|).$ (1)
Taking a derivative in $|S|$ we obtain, that for all $|S|$ such that $d\geq(C-1)\log_{2}|S|+(C-1)/\ln 2$, the right-hand side of (1) is increasing with $|S|$. Consider $|S|\leq d$. Then,
$d-(C-1)\log_{2}(|S|)\geq d-(C-1)\log_{2}d\geq\frac{1.1C\log_{2}C}{\ln 2}>\frac{C}{\ln 2}\,,$
where we used that $d\geq t\geq 5C\log_{2}C$ and $C\geq 2$. Hence, $|S|(d-(C-1)\log_{2}|S|)$ is increasing with $|S|$ for $|S|\leq d$, and for all $S$ with $|S|\leq d$ we have that $e(S,S^{C})\geq d$, as required. For $d\leq|S|\leq 2^{\frac{d-1}{C-1}}$ we have $d-(C-1)\log_{2}|S|\geq 1$, and thus $e(S,S^{C})$ is at least $d$, as required. Thus, suppose that $2^{\frac{d-1}{C-1}}\leq|S|\leq\frac{n}{2}$. Note that $|S|\leq n/2$, which implies that $\log_{C}(n/|S|)\geq\log_{C}2$. Hence, by 1.2,
$e(S,S^{C})$ $\geq\frac{|S|\log_{C}(|S|/2)}{C-1}\geq\frac{|S|\log_{C}2}{C-1}\geq\frac{2^{\frac{d-1}{C-1}}\log_{C}2}{C-1}\geq\frac{2^{\frac{d-1}{C-1}}}{2C\ln C}\,.$ (2)
We claim that for our choice of $t$, (2) is at least $d$. Note that the $\frac{2^{\frac{d-1}{C-1}}}{2C\ln C}-d$ is increasing as $d$ is increasing, and indeed
$\frac{2^{\frac{5C\log_{2}C-1}{C-1}}}{2C\ln C}-5C\log_{2}C>\frac{C^{3}}{2\ln C}-5C\log_{2}C\geq 0,$
@@
The proof of Theorem 1 then follows immediately from Lemma 3.1, since every $d$-regular, $(d-1)$-edge-connected graph has a perfect matching (see [2, Theorem 7, Chapter 18]).
3.1 Threshold for a perfect matching
Recall the definitions of an obstruction and the sets $U,V_{1},W,S$ and $B$ given in Section 1.2, and that $p$ satisfies $(1-p)^{d}=n^{-(1-\epsilon)}$ for a sufficiently small constant $\epsilon\geq 0$. We begin by collecting several lemmas which we will utilise to show that whp there are no obstructions with $u\geq 1$.
3.2 Typical properties of obstructions
We begin with the following Lemma, which will allow us to bound the number of possible obstructions more efficiently.
@@ 3.2 Typical properties of obstructions
Recall that $W$ is a set of $w/2$ edges. We thus have at most $\binom{nd/2}{w/2}\leq(nd)^{w/2}$ ways to choose $W$. As for $S$, it has $\ell_{2}$ components. Let us denote the sizes of these components by $s_{1},\ldots,s_{\ell_{2}}$, where we have that $3\leq s_{i}\leq d^{2}$ for every $i\in[\ell_{2}]$. Thus there are at most $d^{2\ell_{2}}$ ways to choose $s_{1},\ldots,s_{\ell_{2}}$. We then have at most $\binom{n}{\ell_{2}}\leq n^{\ell_{2}}$ ways to choose roots for some spanning trees of these components. Thus by Lemma 1.3, the number of ways to choose $S$ is at most $d^{2\ell_{2}}n^{\ell_{2}}\prod_{i=1}^{\ell_{2}}(ed)^{s_{i}-1}\leq n^{\ell_{2}}(ed)^{2s}$. Finally, $B$ has $\ell_{3}$ components, and here we use the crude bound of at most $n^{\ell_{3}}$ ways to choose their sizes. Then, similarly to the above, by Lemma 1.3 there are at most $n^{2\ell_{3}}(ed)^{b}$ ways to choose $B$.
By Lemma 3.2, given $W\cup S\cup B$ and $u$ there are at most $2$ minimal obstructions. Thus, given $u,w,\ell_{2},s,\ell_{3},$ and $b$, the number of minimal obstructions is at most $2n^{w/2+\ell_{2}+2\ell_{3}}(ed)^{w/2+2s+b}$. ∎
Noting that there are at most $du$ edges touching $U$ in $G$, let us bound from below the number of edges leaving the components $K_{i}$ for $i\in[\ell]$. To that end, for $m\in[n/2]$, let
$f(m)=\max\left\{m(d-(C-1)\log_{2}m),\frac{m\log_{C}(n/m)}{C-1}\right\},$ (3)
where for $m\in[n/2+1,n]$ we set $f(m)=f(n-m)$.
Lemma 3.4.
Suppose $U\subseteq V$ is an obstruction. Then
$\sum_{i=1}^{\ell}e(K_{i},K_{i}^{C})\geq d(\ell-1)+f(k-\ell+1),$
where $f$ is defined according to (3).
Proof.
By Theorem 1.2, we have that $\sum_{i=1}^{\ell}e(K_{i},K_{i}^{C})\geq\sum_{i=1}^{\ell}f(k_{i})$. Further,
$\sum_{i=1}^{\ell}f(k_{i})=\sum_{i=1}^{\ell}\max\left\{k_{i}(d-(C-1)\log_{2}k_{i}),\frac{k_{i}\log_{C}(n/k_{i})}{C-1}\right\}.$
We claim that the minimum of the above sum is obtained with $k_{i}=1$ for all $i\in[\ell-1]$, and $k_{\ell}=k-(\ell-1)$. To that end, consider the discrete functions $g_{1}(x)=x(d-(C-1)\log_{2}x)$ and $g_{2}(x)=\frac{x\log_{C}(n/x)}{C-1}$, supported on $\{1,\ldots,n/2\}$, noting that these functions are concave. Let $g=\max\{g_{1},g_{2}\}$, supported on $\{1,\ldots,n/2\}$. It suffices to show that $g$ is concave. Indeed, considering $\{1,\ldots,n/2\}^{\ell}$, we then have a function $k_{1},\ldots,k_{\ell}\to\sum_{i=1}^{\ell}g(k_{i})$, which is the sum of concave functions, and thus concave in itself, and is defined on a convex set and thus attains its minimum at the extremal points, that is, when $k_{i}=1$ for all $i\in[\ell-1]$ and $k_{\ell}=k-(\ell-1)$. To show that $g$ is concave, it suffices to show that its second derivative is decreasing, that is, that $q(m)=g(m)-g(m-1)$ is decreasing. Observe that there is some minimal integer $m_{0}$ such that for all $m\geq m_{0}$, $g_{2}(m)\geq g_{1}(m)$, and for all $m^{\prime}<m_{0}$, $g_{2}(m^{\prime})\leq g_{1}(m^{\prime})$. Thus, $q$ is decreasing on $\{2,\ldots,m_{0}-1\}$ and on $\{m_{0}+1,\ldots,n/2\}$. All that is then left is to verify that $q(m_{0}+1)\leq q(m_{0})\leq q(m_{0}-1)$, which follows from tedious yet standard calculations, using $t\leq d\leq(C-1)t$ and $2^{t}\leq n\leq C^{t}$. ∎
@@
We finish this section by showing that given an obstruction with $u\geq 2^{\frac{d}{10C}}$, whp $U$ lies in a set which is not ‘too disconnected’, and whose size is proportional to that of $U$. First, we require the following fairly simple claim.

### v3

@@ 1.2 Proofs outline
For larger values of $u$, the aforementioned probabilistic bound no longer suffices. Thus, the final ingredient, similar in spirit to the approach of Bollobás in [4], will be Lemma 3.5, which allows to efficiently bound the number of choices for sets $B$ that have ‘bad expansion’, that is, the number of edges in $E(B,B^{C})$ is small. We note that this is the only place where we use the structure of $G$ as a product graph. Indeed, sets $B$ with bad expansion leave a large ‘fingerprint’ on some coordinates of the product, in particular, if $B$ intersects non-trivially with a projection $G(I)$ of $G$ onto some coordinates $I\subseteq[t]$, then $G(I)$ spans many edges of $E(B,B^{C})$.
@@
Finally, let us briefly comment on the difference between the proof here and the proof for the hypercube, given in [4]. Since the hypercube is a bipartite graph, it suffices there to apply Hall’s theorem, and in particular, one does not need to consider the entirety of the giant component, but only sets of size at most $\frac{n}{2}$ which expand badly. Hence, Harper’s isoperimetric inequality for the hypercube, together with a fingerprint argument given by Bollobás [4], allows one to show that whp there are no obstructions (bar isolated vertices) to a perfect matching. Here, since $G$ is not necessarily bipartite, one needs to consider Tutte-like obstructions, and in particular sets encompassing the entirety of the giant component. This, in turn, requires a much more delicate treatment, and in particular, to consider the structure of obstructions, and the more efficient enumerations given by Lemmas 3.3 and Lemma 3.4 as well as the fact that for a certain range of parameters an obstruction $U$ is typically contained in a well-connected set of the same order (see Lemma 3.14 for a precise statement). This part of the analysis is the key novelty of this work.
@@ 2 Connectivity
Throughout this section, let $\omega=\omega(d)$ be a function tending arbitrarily slowly to infinity and let $p$ be such that $(1-p)^{d}=\omega/n$. We begin by showing that isolated vertices in $G_{p}$ are typically at distance at least two in $G$.
@@
Fix an edge $\{u,v\}\in E$. Since $G$ is $d$-regular, the total number of edges meeting $u$ and $v$ is 2d12d-1. Thus, the probability that $u$ and $v$ are isolated vertices in $G_{p}$ is at most
$(1-p)^{2d-1}\leq(1-p)^{-1}\left(\frac{\omega}{n}\right)^{2},$
where we used that $(1-p)^{d}=\omega/n$. There are $\frac{dn}{2}$ edges to consider. Thus, by the union bound, the probability that two isolated vertices in $G_{p}$ are at distance one in $G$ is at most $\frac{dn}{2}\left(\frac{\omega}{n}\right)^{2}(1-p)^{-1}=o(1)$. ∎
Let us further show the following ‘gap’ statement, which is an almost immediate corollary of Theorem 1.3 and Lemma 1.4. We note that $d^{20}$ in the following lemma is chosen rather arbitrarily, and a much tighter bound can be obtained.
@@
Fix $k\in[2,d^{20}]$. Let us estimate the probability that there is a connected component $K$ of order $k$ in $G_{p}$. By Theorem 1.3, we have that $e(K,K^{C})\geq k(d-(C-1)\log_{C}k)\geq\frac{9kd}{10}$, where we used the fact that $k\leq d^{20}$. Let $\mathcal{T}_{k}$ be the set of trees of order $k$, and let $\mathcal{A}_{k}$ be the event that there exists a connected component of order $k$ in $G_{p}$. Thus, by the union bound and by Lemma 1.4,
$\mathbb{P}\left[\mathcal{A}_{k}\right]$ $\leq\sum_{T\in\mathcal{T}_{k}}(1-p)^{e(T,T^{C})}\leq n(ed)^{k-1}(1-p)^{\frac{9kd}{10}}\leq n\left(\frac{\omega}{n}\right)^{9k/10}(ed)^{k-1}\leq n^{1-4k/5}.$
@@
Whp, there is a unique giant component in $G_{p}$ whose order is no(n)n-o(n), and all the other components in $G_{p}$ are isolated vertices. Furthermore, every two isolated vertices are at distance at least two from each other.
@@
Let $p_{2}=\frac{1}{d^{2}}$, and let $p_{1}$ be such that $(1-p_{1})(1-p_{2})=1-p$. Note that there exists a function $\omega^{\prime}$ tending to infinity such that $(1-p_{1})^{d}=\omega^{\prime}/n$, and thus we may apply Lemmas 2.1 and 2.2 on $G_{p_{1}}$. Furthermore, observe that $G_{p}$ has the same distribution as $G_{p_{1}}\cup G_{p_{2}}$.
We begin by percolating with probability $p_{1}$. Note that by Lemma 2.2, whp in $G_{p_{1}}$ there are only isolated vertices and components of order at least $d^{20}$. We begin by showing that after sprinkling with $p_{2}$, typically all the components of order at least $d^{20}$ merge. Let $L$ be the set of vertices in components of order at least $d^{20}$ in $G_{p_{1}}$. Let $A\sqcup B=L$ be a partition of $L$ which respects the components of $G_{p_{1}}$, that is, every component is contained either in $A$ or in $B$. We may assume that $|A|\leq|B|$ and let $a\coloneqq|A|$.
Let $A^{\prime}\coloneqq\left(A\cup N_{G}(A)\right)\setminus B$, where $N_{G}(X)$ is the external neighbourhood of $X$ in $G$, and let $B^{\prime}\coloneqq\left(B\cup N_{G}(B)\right)\setminus A^{\prime}$. By Lemma 2.1, whp every two isolated vertices in $G_{p_{1}}$ are at distance at least two from each other in $G$. We continue assuming this holds deterministically. Thus, every isolated vertex in $G_{p_{1}}$ is in the neighbourhood of $A$ or $B$, and we have that $V=A^{\prime}\sqcup B^{\prime}$, that is, $B^{\prime}=V\setminus A^{\prime}$. By Theorem 1.3, we have that $e(A^{\prime},B^{\prime})\geq\frac{a\cdot e\cdot\ln(n/a)}{C}\geq\frac{a\cdot e\cdot\ln 2}{C}$, since $a\leq\frac{n}{2}$. We can extend these edges to $\frac{a}{C}$ paths of length at most $3$ between $A$ and $B$. Every two such paths share at most two edges. Thus, we can trim the set of $\frac{a}{C}$ paths to a set of $\frac{a}{Cd^{2}}$ edge-disjoint paths between $A$ and $B$.
We now sprinkle with probability $p_{2}$. The probability none of these paths are in $G_{p_{2}}$ is thus at most $(1-p_{2}^{3})^{\frac{a}{Cd^{2}}}\leq\exp\left\{-\frac{a}{d^{9}}\right\}$. For a union bound note that there are at most $\binom{n/d^{20}}{a/d^{20}}$ possibilities to choose the partition of $L$. Thus, the probability that there is a partition of $L$ such that there is no path between two components of order at least $d^{20}$ is at most
@@
Note that by Lemma 2.1, whp every two isolated vertices in $G_{p_{1}}$ are not connected by an edge of $G$. Hence, adding any edge touching an isolated vertex connects it to a component whose order is at least $d^{20}$ in $G_{p_{1}}$, and these components all merge whp. Hence, whp, there exists a unique connected component in $G_{p_{1}}$ whose order is at least $d^{20}$, and all the other components are isolated vertices, whose distance in $G$ is at least two. Finally, let $Y$ be the random variable counting the number of isolated vertices in $G_{p}$. Then, $\mathbb{E}[Y]=n(1-p)^{d}\leq\omega$. Thus, by Markov’s inequality, whp $Y\leq\omega^{2}=o(n)$, completing the proof. ∎
@@ 3 Perfect matching
This section is structured as follows. In Subsection 3.1 we prove Theorem 1. Then, in Subsection 3.2 we prove several estimates on the number of obstructions and related sets. Finally, in Subsection 3.3 we prove Theorem 3b. Throughout the section, we assume divisibility by two whenever is necessary, noting that minor modifications will allow us to argue for the existence of nearly-perfect matching otherwise.
3.1 Proof of Theorem 1
We begin with the proof of Theorem 1. Let us first show that, by Theorem 1.3, for ‘high-enough’ dimension of $G$ every edge-cut of $G$ has at least $d$ edges.
@@ 3.1 Proof of Theorem 1
Suppose that $t\geq 5C$. Then, we have that $G$ is $d$-edge-connected.
@@
It suffices to show that $e(S,S^{C})\geq d$ for all $S\subseteq V$ with $1\leq|S|\leq n-1$, and in fact by symmetry, for $S$ of size at most $\frac{n}{2}$. By 1.3,
$e(S,S^{C})\geq|S|(d-(C-1)\log_{C}|S|).$ (1)
Taking a derivative in $|S|$ we obtain, that for all $|S|$ such that $d\geq(C-1)\log_{C}|S|+(C-1)/\ln 2$, the right-hand side of (1) is increasing with $|S|$. Consider $|S|\leq d$. Then,
$d-(C-1)\log_{C}(|S|)\geq d-(C-1)\log_{C}d>2C\,$,
where we used that $d\geq t\geq 5C$ and $C\geq 2$. Hence, $|S|(d-(C-1)\log_{2}|S|)$ is increasing with $|S|$ for $|S|\leq d$, and for all $S$ with $|S|\leq d$ we have that $e(S,S^{C})\geq d$, as required. For $d\leq|S|\leq C^{\frac{d-1}{C-1}}$ we have $d-(C-1)\log_{C}|S|\geq 1$, and thus $e(S,S^{C})$ is at least $d$, as required. Thus, suppose that $C^{\frac{d-1}{C-1}}\leq|S|\leq\frac{n}{2}$. Note that $|S|\leq n/2$, which implies that $\ln(n/|S|)\geq\ln 2$. Hence, by 1.3,
$e(S,S^{C})$ $\geq\frac{|S|\cdot e\cdot\ln(n/|S|)}{C}\geq\frac{|S|}{C}\geq\frac{C^{\frac{d-1}{C-1}}}{C}=C^{\frac{d-1}{C-1}-1}\,.$ (2)
We claim that for our choice of $t$, (2) is at least $d$. Note that the $C^{\frac{d-1}{C-1}}-1-d$ is increasing as $d$ is increasing, and indeed
$C^{\frac{5C-1}{C-1}-1}-5C>C^{3}-5C\geq 0,$
@@
The proof of Theorem 1 then follows immediately from Lemma 3.1, since every $d$-regular, (d1)(d-1)-edge-connected graph has a perfect matching (see [2, Chapter 18, Theorem 7]).
3.2 Properties of obstructions
Recall from Section 1.2 the definition of a minimal obstruction, further recall the sets $U,V_{1},W,S$, and $B$, and $\ell,\ell_{1},\ell_{2},\ell_{3}$. In this Section, we obtain several deterministic lemmas, which will allow us to count the number of minimal obstructions more efficiently.
The first lemma is a rather simple (yet, as we will see in Lemma 3.13, quite powerful) observation.
Lemma 3.2.
Let $U$ be a minimal obstruction. Then any vertex $v\in U$ has neighbours in at least three components of $G_{p}[V_{1}\cup S\cup B]$.
Proof.
Suppose towards contradiction that there is a minimal obstruction $U$ of size $u$ with $\ell>u$ components in $G_{p}[V_{1}\cup S\cup B]$ containing a vertex $v$ whose neighbourhood intersects at most two components of $G_{p}[V_{1}\cup S\cup B]$.
Consider $U^{\prime}=U\setminus\{v\}$. Clearly, $|U^{\prime}|=u-1$. Let us denote by $V_{1}^{\prime},S^{\prime},B^{\prime}$ the sets containing (respectively) components of size exactly one, between $3$ and $\frac{n}{d^{C^{3}/p}}$, and larger than $\frac{n}{d^{C^{3}/p}}$ in $G_{p}[V\setminus U^{\prime}]$. Observe that since $v$ is connected to at most two components in $G_{p}[V_{1}\cup S\cup B]$, there are at least $\ell-1>u-1=|U^{\prime}|$ components in $G_{p}[V_{1}^{\prime}\cup S^{\prime}\cup B^{\prime}]$. But then $U^{\prime}$ is an obstruction, contradicting the minimality of $U$. ∎
The following observation will be used in the subsequent lemma, Lemma 3.4, to count obstructions efficiently by choosing $W,S$ and $B$ (instead of $U$).
@@ 3.2 Properties of obstructions
Recall that $W$ is a set of $w/2$ edges. We thus have at most $\binom{nd/2}{w/2}\leq(nd)^{w/2}$ ways to choose $W$. As for $S$, it has $\ell_{2}$ components. Let us denote the sizes of these components by $s_{1},\ldots,s_{\ell_{2}}$, where we have that $3\leq s_{i}\leq\frac{n}{d^{C^{3}/p}}$ for every $i\in[\ell_{2}]$. Thus there are at most $(\frac{n}{d^{C^{5}/p}})^{\ell_{2}}\leq n^{\ell_{2}}$ ways to choose $s_{1},\ldots,s_{\ell_{2}}$. We then have at most $\binom{n}{\ell_{2}}\leq n^{\ell_{2}}$ ways to choose roots for some spanning trees of these components. Thus by Lemma 1.4, the number of ways to choose $S$ is at most $n^{2\ell_{2}}\prod_{i=1}^{\ell_{2}}(ed)^{s_{i}-1}\leq n^{2\ell_{2}}(ed)^{s}$. Finally, $B$ has $\ell_{3}$ components, and there are at most $n^{\ell_{3}}$ ways to choose their sizes. Then, similarly to the above, by Lemma 1.4 there are at most $n^{2\ell_{3}}(ed)^{b}$ ways to choose $B$.
By Lemma 3.3, given $W\cup S\cup B$ and $u$ there are at most $2$ minimal obstructions. Thus, given $u,w,\ell_{2},s,\ell_{3},$ and $b$, the number of minimal obstructions is at most $2n^{w/2+2\ell_{2}+2\ell_{3}}(ed)^{w/2+s+b}$. ∎
@@
3.3 Proof of Theorem 3b
Recall from Section 1.2 the definition of a minimal obstruction and of a non-trivial minimal obstruction. Further, recall the sets $U,V_{1},W,S$, and $B$, and $\ell,\ell_{1},\ell_{2},\ell_{3}$. Recall that $\omega$ is a function tending arbitrarily slowly to infinity and that $p$ is such that $(1-p)^{d}=\omega/n$.
Before turning to the proof itself, we require some preparation.
For $k\in\{1,\ldots,n\}$, we define
$f(k)=\min_{A\subseteq V(G),|A|=k}e(A,A^{C}).$ (3)
We note that f(k)=f(nk)f(k)=f(n-k), and we may thus restrict our attention to $k\in\{1,\ldots,n/2\}$. Further, by Theorem 1.3,
$f(k)\geq f^{*}(k)\coloneqq\max\left\{k(d-(C-1)\log_{C}k),k\cdot\frac{e}{C}\cdot\ln\left(\frac{n}{k}\right)\right\}.$
Remark 3.6.
We note that, while in general the maximum of two concave functions is not necessarily concave, $f^{*}(k)$ is indeed concave in the domain $\{1,\ldots,n/2\}$ (see, e.g., [12, Section 3.4]). We further note that $f^{*}$ is increasing in $\{1,\ldots,n/e\}$, and decreasing afterwards. Further, $f^{*}(n/2)\geq n/C$.
Given an obstruction $U$, let $W_{1},\dots,W_{w/2}$ denote the components of $G_{p}[W]$, $S_{1},\dots,S_{\ell_{2}}$ denote the components of $G_{p}[S]$ and $B_{1},\dots,B_{\ell_{3}}$ denote the components of $G_{p}[B]$ and set
$\hat{e}\coloneqq d\ell_{1}+\sum_{i=1}^{w/2}e(W_{i},W_{i}^{C})+\sum_{i=1}^{\ell_{2}}e(S_{i},S_{i}^{C})+\sum_{i=1}^{\ell_{3}}e(B_{i},B_{i}^{C}).$
We will utilise the following two edge-counting claims. The first concerns $G_{p}[S]$.